Military Draft. Determine the transfer. ($millions)
D: Q = 850
S: P = 200 + 4Q
Draft: P = 50
1
Expert's answer
2016-05-24T12:43:03-0400
Military Draft.
D: Q = 850,
S: P = 200 + 4Q.
The equilibrium price P = 200 + 4*850 = $3,600.
If the amount of draft is 50, then the transfer is 50*3,600 = $180,000.